OpenStudy (anonymous):

d=13cm h=24cm how do you find the volume in terms of pi

OpenStudy (ericanoel912):

What shape is it?

OpenStudy (midhun.madhu1987):

Cone or Cylinder?

OpenStudy (anonymous):

Vcone= \[\pi \times r ^{2}\times \frac{ h }{ 3 }\] Vcylinder= \[\pi \times r ^{2}\times h\]

OpenStudy (anonymous):

volume = area * h and area = [3.14*(d)^2]/4 area = [3.14*(13^2)]/4 = cm^2 volume = 24 * area =cm^3

OpenStudy (anonymous):

A cylinder